OpenStudy (anonymous):

The daily rainfall during two April weeks was: 1, 0.5, 0.3, 0, 0 ,0, 1.2, 3, 0, 1.1, 0.7, 2, 1.3, 2 inches. What was the median during the two week period?

OpenStudy (anonymous):

First arrange the data in ascending order (i.e.) from smallest to biggest: 0, 0, 0, 0, 0.3, 0.5, 0.7, 1, 1.1, 1.2, 1.3, 2, 2, 3 Now,the median is the (n+1)/2th term , here n=14 So, (n+1)/2 = 15/2 = 7.5 7.5th term is the average of 7th and 8th term i.e. average of 0.7 and 1, which is 0.85 So median is 0.85

OpenStudy (anonymous):

Thank youu